OpenStudy (dimpalk104):

please help me with conversion factors If a person drinks an average of 1.45 L of Diet Coke every day since 1987, how many gallons of Diet Coke has he/she consumed in the past 17 years?

OpenStudy (anonymous):

Without showing you numbers can you figure it out. \[\frac{years }{ 1 }\times \frac{ days }{ year }\times \frac{ Liters }{ days }\times \frac{ gallons }{ Liters }\]

OpenStudy (dimpalk104):

no not really because aren't you supposed to start with 1.45/ 1 day?? @Shalante

OpenStudy (anonymous):

No, you would use the given variable in the end of the last sentence to begin with (years)

OpenStudy (anonymous):

Tell me how your conversion would look in liters? If we want to find gallons, we know that liter can be convert to gallon If you drink 1.45 L a day and we know how much days are in a year, then we can figure it out.

OpenStudy (anonymous):

It is like a step. The unit you are trying to find should be in the last conversion fraction.

OpenStudy (dimpalk104):

i got 2377 as my answer @Shalante

OpenStudy (anonymous):

Let me see.

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

I gtg 365 years*365days/1 year * 1.45L/day* 0.26 gallon/1Liters

OpenStudy (anonymous):

17 years oops.

OpenStudy (anonymous):

@Shalante Can't you do the following: \[\frac{ 1.45L }{ day }(\frac{ 0.264172 gallons }{ L }) (\frac{ 365days }{ year })(\frac{ 17 yrs }{ 1 })\] This would equal 2376.82 as in @Dimpalk104 answer

OpenStudy (anonymous):

Starting with years is easier for me since you would cancel the unit in the numerator every time. It looks more chronological too. Your way seems to be a bit more confusing as it goes from gallons in numerator to years in denominator.

OpenStudy (anonymous):

Fine, but either way you would get the same answer @Dimpalk104 got 2377 as the answer. I'm wondering why you responded nope

OpenStudy (anonymous):

Thought he/she put something else because I saw the answer and responded a little late. Maybe he/she deleted the answer. But that's still not the answer because it is not in 3 sig figs.

OpenStudy (anonymous):

Your way of starting out and ending the conversion has its flaws that would not be helpful for the more different and complex ones.